Find the solution of the initial value problem y(1) = 46 y = sqrt(e^(-1/x+2)+5) Your answer should be a function of x.
Answer #1

7 5 9 = 1 y. [yetlero] (088 use the method of variable seprable 3 (no S 124 1 day = S t adik let y2_5=47(24 dy = (se? du -zti92-5= e(1-1/4) y² = e (1 - 1 - - y= Venta + 5 no y = sqrt (e**)+5))